# How Bright in Gravitational Waves are Millisecond Pulsars for the Galactic Center GeV Gamma-Ray Excess? A Systematic Study

### v1

@@ VI Results and discussion
The relative placement of the scattered points with respect to the solid sensitivity curves  [^5] The sensitivity curves for LVK O4 are adopted from Ref. Abac et al. (2025). The PSD data files for CE (with the baseline 40 km detector) and ET (with option ET-B) are available for download at https://dcc.cosmicexplorer.org/CE-T2000017/public and https://www.et-gw.eu/index.php/etsensitivities, respectively. illustrates the prospects of detecting those individual MSPs in the coherent strategy, while in the incoherent strategy, detectability is assessed by comparing the green dashed line with the dashed sensitivity curves. Overall, for the coherent strategy, the predicted $h_{0}$ values are below the reach of current detectors Aasi et al. (2015); Acernese et al. (2015); Akutsu et al. (2019) but can be detectable by next-generation detectors like CE and ET Evans et al. (2023); Punturo et al. (2010). For the incoherent strategy, only the theoretical maximum from Pop. I marginally exceeds the sensitivity of CE with a one-year observation time.
@@
Finally, the bottom row of Fig. 3 shows the theoretical maximum case, where $\eta=100\%$, i.e., all spin-down energy is converted into GW emission. In this maximum scenario, both $h_{0}$ and $h_{\rm eff}$ are a factor of 10 larger than in the $\eta=1\%$ case in the third row, representing the absolute upper bound on the GW emission from MSPs. Remarkably, LVK O4 can observe a fraction of the MSPs in the high-$h_{0}$ tail, implying that Advanced LIGO and Virgo at design sensitivity can potentially detect some individual sources Capote et al. (2025); Abac et al. (2025). Conversely, a null detection in O4 can place meaningful constraints on such optimistic scenarios, carrying important implications for the maximum possible ellipticities of MSPs and their role in explaining the GCE.
@@ VI.1 Compare with previous work
Our detectability is a few orders of magnitude lower compared with Ref. Miller and Zhao (2023) by Miller & Zhao. The main reason is that Ref. Miller and Zhao (2023) included PSR J0537-6910 in their ATNF-catalog-drive MSP population model, which dramatically boosted their detectability estimate.
PSR J0537-6910 has quite different properties from MSPs. Although its period is 16.2 ms, only slightly larger than that of typical MSPs (1–10 ms), its period derivative is $\dot{P}=5.2\times 10^{-14}$, many orders of magnitude larger than those of typical MSPs ($10^{-21}$$10^{-19}$). As a result, the spin-down power is $\dot{E}\simeq 5\times 10^{38}$ erg/s, enormously higher than typical MSPs ($10^{33}$$10^{35}$ erg/s). The corresponding spin-down age is only $\simeq 4.9\times 10^{3}$ yr, and the derived surface magnetic strength is $\simeq 9.3\times 10^{11}$ Gs, both consistent with a young pulsar rather than an MSP ($\sim 1$ Gyr and $\sim 10^{8}$$10^{9}$ Gs). Thus, Ref. Miller and Zhao (2023) derived an ellipticity of $\simeq 10^{-6}$ for PSR J0537-6910, which is a strong outlier in their ellipticity distribution (see their Figs. 4 and 5), and this drives their detectability by orders of magnitude. We do not include PSR J0537-6910 because its above properties indicate that it is a young pulsar rather than an MSP. Moreover, PSR J0537-6910 is located in the Large Magellanic Cloud (LMC) rather than in the Milky Way  [^6] Despite not being a member of the MSP population involved in our Pop. I samples, it remains a promising candidate for targeted GW searches Fesik and Papa (2020); Abbott et al. (2021b, c); Abac et al. (2025)..
@@
Next, compared with Ref. Bartel and Profumo (2024) by Bartel & Profumo, our detectability in the coherent strategy is much better. The main reason is that the sensitivity curve used in Ref. Bartel and Profumo (2024) appears to be taken directly from the detector PSD, whereas in our calculation, we use Eq. (23), which incorporates the exposure time for continuous GWs.
@@ VII Conclusion
The GCE has long been considered a leading candidate signal for particle dark matter annihilation. However, an unresolved population of MSPs in the Galactic bulge offers the other compelling astrophysical explanation. Distinguishing between these two interpretations is challenging in electromagnetic observations due to source confusion, pulse broadening, and extinction. Gravitational waves, by contrast, provide a dust- and confusion-free probe: a steadily rotating, non-axisymmetric MSP emits a nearly monochromatic GW at about twice its rotational frequency, within the sensitive band of ground-based GW detectors, with amplitude mainly determined by its ellipticity.
@@
Our main result, on the GW strain amplitudes and their detectability, is shown in Fig. 3. Overall, in the coherent strategy, the predicted strain amplitudes, $h_{0}$, lie below the reach of current LVK sensitivities, but can be detectable by next-generation detectors such as CE and ET. In contrast, in the incoherent strategy, only the theoretical maximum from Pop. I marginally exceeds the sensitivity of CE. In addition, it is worth noting that the true detectability can be more optimistic than suggested by our baseline forecasts, since the strain amplitudes are primarily governed by two factors: (1) the total number of bulge MSPs, $N_{\rm MSP}$, which controls the extent of the high-$h_{0}$ tail and the summed $h_{\rm eff}$’s normalization, and (2) the physical mechanisms that determine the ellipticity in each origin, which we adopt conservative assumptions. Finally, in the theoretical maximum case with $\eta=100\%$, where all spin-down energy is converted into GW emission, the LVK O4 can potentially detect a fraction of MSPs. Conversely, a null detection in LVK O4 can constrain such optimistic scenarios, with important implications for the maximum sustainable ellipticities of MSPs and their viability in explaining the GCE.
@@
In summary, GW observations offer a unique and complementary window into the Galactic bulge MSP population, overcoming the severe challenges faced in electromagnetic searches. This provides a crucial way to test the MSP interpretation of the GCE, which is the alternative (astrophysical) explanation to the particle dark matter interpretation. Looking ahead, the next generation of GW interferometers will either reveal a long-sought MSP population near the Galactic center or place stringent limits on their ellipticities and abundance. Either outcome will carry profound implications for the origin of the GCE, the role of recycled neutron stars in shaping Galactic high-energy phenomena, and the broader quest to disentangle conventional astrophysical sources from potential dark-matter and other new-fundamental-physics signatures.
@@ Acknowledgments
We are grateful for helpful discussions with Kayla Bartel, John Beacom, Emanuele Berti, Marc Kamionkowski, Gordan Krnjaic, Andrew Miller, Benjamin Owen, Luca Reali, Yue Zhao, and Haowen Zhong. B. Z. is supported by Fermi Forward Discovery Group, LLC under Contract No. 89243024CSC000002 with the U.S. Department of Energy, Office of Science, Office of High Energy Physics. X.H. is supported by the National Key Research and Development Program of China (2022YFF0503304), the National Natural Science Foundation of China (No. 12322302), and the Project for Young Scientists in Basic Research of Chinese Academy of Sciences (No. YSBR-061).

### v2

@@ VI Results and discussion
The relative placement of the scattered points with respect to the solid sensitivity curves  [^10] The sensitivity curves for LVK O4 are adopted from Ref. Abac et al. (2025a). The PSD data files for CE (with the baseline 40 km detector) and ET (with 20 km arms) Branchesi et al. (2023); Abac et al. (2025b) are available for download at https://dcc.cosmicexplorer.org/CE-T2000017/public and https://apps.et-gw.eu/tds/ql/?c=16492, respectively. illustrates the prospects of detecting those individual MSPs in the coherent strategy, while in the incoherent strategy, detectability is assessed by comparing the green dashed line with the dashed sensitivity curves. Overall, for the coherent strategy, the predicted $h_{0}$ values are below the reach of current detectors Aasi et al. (2015); Acernese et al. (2015); Akutsu et al. (2019) but can be detectable by next-generation detectors like CE and ET Evans et al. (2023); Punturo et al. (2010). For the incoherent strategy, only the theoretical maximum from Pop. I marginally exceeds the sensitivity of CE and ET with a one-year observation time.
@@
Finally, the bottom row of Fig. 3 shows the theoretical maximum case, where $\eta=100\%$, i.e., all spin-down energy is converted into GW emission. In this maximum scenario, both $h_{0}$ and $h_{\rm eff}$ are a factor of 10 larger than in the $\eta=1\%$ case in the third row, representing the absolute upper bound on the GW emission from MSPs. Remarkably, LVK O4 can observe a fraction of the MSPs in the high-$h_{0}$ tail, implying that Advanced LIGO and Virgo at design sensitivity can potentially detect some individual sources Capote et al. (2025); Abac et al. (2025a). Conversely, a null detection in O4 can place meaningful constraints on such optimistic scenarios, carrying important implications for the maximum possible ellipticities of MSPs and their role in explaining the GCE.
@@ VI.1 Compare with previous work
Our detectability is a few orders of magnitude lower compared with Ref. Miller and Zhao (2023) by Miller & Zhao Mil . The main reason is that Ref. Miller and Zhao (2023) included PSR J0537-6910 in their ATNF-catalog-drive MSP population model, which dramatically boosted their detectability estimate.
PSR J0537-6910 has quite different properties from MSPs. Although its period is 16.2 ms, only slightly larger than that of typical MSPs (1–10 ms), its period derivative is $\dot{P}=5.2\times 10^{-14}$, many orders of magnitude larger than those of typical MSPs ($10^{-21}$$10^{-19}$). As a result, the spin-down power is $\dot{E}\simeq 5\times 10^{38}$ erg/s, enormously higher than typical MSPs ($10^{33}$$10^{35}$ erg/s). The corresponding spin-down age is only $\simeq 4.9\times 10^{3}$ yr, and the derived surface magnetic strength is $\simeq 9.3\times 10^{11}$ Gs, both consistent with a young pulsar rather than an MSP ($\sim 1$ Gyr and $\sim 10^{8}$$10^{9}$ Gs). Thus, Ref. Miller and Zhao (2023) derived an ellipticity of $\simeq 10^{-6}$ for PSR J0537-6910, which is a strong outlier in their ellipticity distribution (see their Figs. 4 and 5), and this drives their detectability by orders of magnitude. We do not include PSR J0537-6910 because its above properties indicate that it is a young pulsar rather than an MSP. Moreover, PSR J0537-6910 is located in the Large Magellanic Cloud (LMC) rather than in the Milky Way  [^11] Despite not being a member of the MSP population involved in our Pop. I samples, it remains a promising candidate for targeted GW searches Fesik and Papa (2020); Abbott et al. (2021c, d); Abac et al. (2025a)..
@@
Next, compared with Ref. Bartel and Profumo (2024) by Bartel & Profumo Bar , our detectability in the coherent strategy is much better. The main reason is that the sensitivity curve used in Ref. Bartel and Profumo (2024) appears to be taken directly from the detector PSD, whereas in our calculation, we use Eq. (23), which incorporates the exposure time for continuous GWs.
@@ VII Conclusion
The GCE has long been considered a leading candidate signal for particle DM annihilation. However, an unresolved population of MSPs in the Galactic bulge offers the other compelling astrophysical explanation. Distinguishing between these two interpretations is challenging in electromagnetic observations due to source confusion, pulse broadening, and extinction. Gravitational waves, by contrast, provide a dust- and confusion-free probe: a steadily rotating, non-axisymmetric MSP emits a nearly monochromatic GW at about twice its rotational frequency, within the sensitive band of ground-based GW detectors, with amplitude mainly determined by its ellipticity.
@@
Our main result, on the GW strain amplitudes and their detectability, is shown in Fig. 3. Overall, in the coherent strategy, the predicted strain amplitudes, $h_{0}$, lie below the reach of current LVK sensitivities, but can be detectable by next-generation detectors such as CE and ET. In contrast, in the incoherent strategy, only the theoretical maximum from Pop. I marginally exceeds the sensitivity of CE and ET. In addition, it is worth noting that the true detectability can be more optimistic than suggested by our baseline forecasts, since the strain amplitudes are primarily governed by two factors: (1) the total number of bulge MSPs, $N_{\rm MSP}$, which controls the extent of the high-$h_{0}$ tail and the summed $h_{\rm eff}$’s normalization, and (2) the physical mechanisms that determine the ellipticity in each origin, which we adopt conservative assumptions. Finally, in the theoretical maximum case with $\eta=100\%$, where all spin-down energy is converted into GW emission, the LVK O4 can potentially detect a fraction of MSPs. Conversely, a null detection in LVK O4 can constrain such optimistic scenarios, with important implications for the maximum sustainable ellipticities of MSPs and their viability in explaining the GCE.
@@
In summary, GW observations offer a unique and complementary window into the Galactic bulge MSP population, overcoming the severe challenges faced in electromagnetic searches. This provides a crucial way to test the MSP interpretation of the GCE, which is the alternative (astrophysical) explanation to the particle DM interpretation. Looking ahead, the next generation of GW interferometers may reveal a long-sought MSP population near the Galactic center or place stringent limits on their ellipticities and abundance. Conversely, a non-detection would constrain the MSP hypothesis, and such a null result would indirectly provide further support for the DM interpretation of the GCE. Either outcome will carry profound implications for the origin of the GCE, the role of recycled neutron stars in shaping Galactic high-energy phenomena, and the broader quest to disentangle conventional astrophysical sources from potential dark-matter and other new-fundamental-physics signatures.
@@ Acknowledgments
We are grateful for helpful discussions with Kayla Bartel, John Beacom, Emanuele Berti, Ilias Cholis, Marc Kamionkowski, Gordan Krnjaic, Andrew Miller, Benjamin Owen, Luca Reali, Nick Rodd, Tracy Slatyer, Yue Zhao, and Haowen Zhong. B. Z. is supported by Fermi Forward Discovery Group, LLC under Contract No. 89243024CSC000002 with the U.S. Department of Energy, Office of Science, Office of High Energy Physics. X.H. is supported by the National Key Research and Development Program of China (2022YFF0503304), the National Natural Science Foundation of China (No. 12322302), and the Project for Young Scientists in Basic Research of Chinese Academy of Sciences (No. YSBR-061).